OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, –4) with a slope of one half. y = 1/2x + 5 y = 1/2x − 3 y = 1/2x − 5 y = 1/2x + 3

OpenStudy (anonymous):

Please help @darkdirl111296

OpenStudy (anonymous):

i saw slope & thats my weakest point sorry i cant

OpenStudy (anonymous):

@darkdirl111296 ok thanks for trying

OpenStudy (anonymous):

np

OpenStudy (danjs):

Given a point on the line, and the slope (x1 , y1) = ( 2 , -4) m = 1/2 point-slope line equation y - y1 = m*(x - x1) put in the m, x1, y1 you can turn it to the answer form if needed